Give an example of a business model in 2 to 5 sentences. (Do not u something similar is fine. Try to challenge yourselft) 4. se the example I used in class -but a. Create a Cost function and a Revenue function that models this business situation. b. Show your calculations to find the profit function here c. Graph your Cost and Revenue functions and label what the intersection represents. d. Discuss when you will start making a profit and how you can determine this from a graph. 5. Given a supply and demand function, respectively Supply p ,2q +1 Demand p+.lg 70 What is the market equilibrium? Explain what this means for Supply and Demand as well as price and quantity. Remember: quantity is the input and price is the output.
Answer #1

SOLUTION:

1) Cost function: C(x) = F + Vx; where F = Fixed cost; V = Variable cost Per unit; x= No of units produced and sold; C = Total cost

 

Revenue function: R(x) = px; where R(x) reflects the revenue from selling

 

Suppose we have Fixed cost = 200,000; Variable cost: $10; Selling price $50 each unit

 

Thus Cost function: C(x) = F + Vx = 200,000 + $10x

Revenue function: R(x) = px = $50x

 

2) Profit Functions: P(x) = R(x) = C(x)

Profit Functions: $50x - 200,000 - $10x = $40x - 200,000

3) Graph enclosed

 

4) GRAPH 400000 300000 20000 100000 Loss C=R Breakeven R C Profit Fixed cost CSR 2000 8000 100000 200000 Breakeven 300000 From break-even point we can determine when the firm starts earning profits. Once the company surpass the break-even price, it can start making a profit. It is clearly indicated in the graph